OpenStudy (anonymous):

Question: 4 = √p - 2 ------- Possible Answers: √6 36 3 6 Four more questions that has to be posted.

Parth (parthkohli):

Is it the square root of whole p-2?

OpenStudy (anonymous):

No, just the √p, separate the - 2, and it's not negative 2.

OpenStudy (angela210793):

@FeoNeo33 will u try urself first please???

Parth (parthkohli):

Simply get the answer by adding 2 to both sides and squaring both sides.

OpenStudy (anonymous):

What's the √p?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 4 + 2 = \sqrt{p} \cancel{- 2 + 2} }\) \(\Large \color{MidnightBlue}{\Rightarrow 6 = \sqrt{p} }\) What do you get by squaring both sides?

OpenStudy (anonymous):

Let's go by the add two to both sides, if, I add 2 to 4,(=6) then squared it, it would come out as 36?

Parth (parthkohli):

Yes. Right. Correct.

OpenStudy (anonymous):

hmm... aight i probably got the rest of the three questions. There's really one that's just... wtf in my eyes. lol thank you very much.